# Learning nuclear cross sections across the chart of nuclides with graph neural networks

### v2

@@ VI Conclusion
In this work, we used variational autoencoders (VAEs) and implicit neural representations (INRs) to learn the latent space of nuclear cross sections, specifically the elastic, inelastic, $(n,2n)$ and $(n,\gamma)$ cross sections in the fast regime. Both types of architecture were trained on the synthetic data from the TENDL nuclear dataset. We then used graph neural networks (GNNs) on the latent representations to learn the relationships between these four types of cross sections. We showed that we could predict cross sections in a rather large test set with remarkably good accuracy. We also outlined how to quantify the impact new measurements or theoretical calculations would have on the GNN predictions.
@@ VII Appendix
In this section, we provide supplementary analyses and ablation studies to support the findings presented in the main paper. These experiments are designed to validate the robustness and generalization capabilities of the proposed framework. We begin by investigating the representation learning phase with the VAE. Following this, we present the results of the downstream GNN prediction task, focusing on consistent performance across different test regions and varying window lengths.
@@ VII.1 Train-Test Performance Trends for VAE
To ensure that our model is not overfitting during the representation learning phase, we analyze the evolution of the loss and PSNR for both the training and testing set during the VAE training. These metrics are presented as a function of epochs in Figure 12. The training and testing loss curves show a steady decrease, converging to similar values by the end of training. Also, the PSNR values for both training and testing increase consistently, improving reconstruction quality over time. The consistent convergence of both metrics indicates minimal overfitting and robust generalization of the learned latent representations.
@@ VII.2 Generalization Analysis in GNN
Performance consistency across test regions
In the downstream prediction task with the GNN, we conduct an analysis to verify the model’s ability to generalize across different regions of the test set. To this end, we perform evaluation on multiple test regions centered on distinct nuclear chart coordinates while keeping the window length fixed (e.g., $\Delta Z=\Delta N=8$). Figure 13 indicates that test PSNR scores across all randomly selected regions are consistent and demonstrates that the model is not biased toward specific regions of the training set. The three curves with the lowest PSNR correspond to regions centered on 214Tl ($Z=81,N=133$), 219At ($Z=85,N=134$) and 239Pa ($Z=91,N=148$). In these three cases, the TENDL library contains overall fewer isotopes of each element; see Fig. 9. As a result, the number of actual nuclei in the test set is much smaller than the 289 nuclei of most other regions. This lack of samples in these regions results in sparse relationships between neighboring nuclei compared to regions with denser neighboring points. This sparsity limits the GNN’s ability to learn comprehensive connections, leading to lower PSNR scores compared to denser test regions.
@@
Figure 13: PSNR values in the test set as a function of epochs for test regions centered at distinct nuclear chart coordinates $[N_{0},Z_{0}]$. All test regions have the same size characterized by $\Delta Z=\Delta N=8$; only the center nucleus changes.
Performance consistency across window length variations
Figure 14 shows the test PSNR scores across different window lengths for the Gated and ARMA approaches. This analysis highlights how the model performs when varying the size of the test region, where increasing the window length reduces the number of training samples available. Such a setup is crucial to evaluate the model’s ability to generalize effectively, as it forces the model to learn and infer from sparse training data. As see in this figure, the model demonstrates robust and consistent performance across a wide range of window lengths, verifying its capability to generalize well even under challenging conditions.
@@
Figure 14: Test PSNR values across varying window sizes for Gated (left) and ARMA (right) approaches, with the center of the test window fixed at $(Z_{0},N_{0})=(42,60)$.
VII.3 Location of latent clusters

### v3

@@ VI Conclusion
In this work, we used variational autoencoders (VAEs) and implicit neural representations (INRs) to learn the latent space of nuclear cross sections, specifically the elastic, inelastic, $(n,2n)$ and $(n,\gamma)$ cross sections in the fast regime. Both types of architecture were trained on the synthetic data from the TENDL nuclear dataset. We then used graph neural networks (GNNs) on the latent representations to predict the missing cross sections across the nuclear chart by leveraging the topological structure. We showed that we could predict cross sections of a rather decent-sized test set with good accuracy using a physics-free methodology: the networks learn from the data, whether the data come from measurements, calculations or evaluations. The fact that this proof-of-concept study used the TENDL dataset was motivated exclusively by the need to have enough data to train the networks.
@@ VII Appendix
This Appendix contains supplementary analyses and ablation studies to support the findings presented in the main paper. We first begin by investigating the representation learning phase, followed by additional analyses of the GNN-based prediction task, including variations in the test region location, test window size, and different training strategies.
@@ VII.1 Train-Test Performance Trends for VAE
To ensure the model does not overfit during the representation learning phase, we analyze the evolution of the loss and PSNR for both the training and testing set during the VAE training. These metrics are presented as a function of epochs in Fig. 9. Both training and testing loss curves steadily decrease and converge to similar values by the end of training. Also, the PSNR values for both training and testing increase consistently, improving reconstruction quality over time. The consistent convergence of both metrics indicates minimal overfitting and robust generalization of the learned latent representations.
@@ VII.2 Generalization Analysis in GNN
Choice of test regions
To investigate the model performance in imputation across different regions of the nuclear nuclear chart, we evaluate the test PSNR scores at eight distinct locations $(Z_{0},N_{0})$ for the center of the test region. These locations vary in their local neighborhood density. For example, test regions centered on medium-mass nuclei such as $(Z_{0},N_{0})=(45,55)$, $(Z_{0},N_{0})=(42,60)$, and $(Z_{0},N_{0})=(57,71)$ are densely surrounded by known neighbors, while test regions centered on heavy nuclei such as $(Z_{0},N_{0})=(81,133)$, $(Z_{0},N_{0})=(85,134)$, or $(Z_{0},N_{0})=(91,148)$ are closer to the boundaries of the domain and are surrounded by fewer neighbors; see Fig.8 for example. As shown in Fig. 10, the imputation performance is strongly correlated with local neighborhood density. Test regions with dense neighboring support yield higher PSNR scores, indicating effective message passage and generalization. In contrast, edge regions with limited topological context yield noticeably lower PSNR, reflecting the increased difficulty of inference. This observation highlights the importance of neighborhood context in graph-based models and the advantages of transductive message-passing when applied within densely populated nuclear regions.
@@
Figure 10: Test PSNR scores for GNN imputation in test regions centered at different $(Z_{0},N_{0})$ across the nuclear chart. The performance trend reflects the influence of local structural support on message-passing effectiveness and prediction quality. All test regions have the same size characterized by $\Delta Z=\Delta N=4$ (i.e., $9\times 9$ test window); only the center nucleus changes.
Choice of test window lengths
To further investigate how the imputation performance is influenced by the spatial extent of the test region, we analyze the GNN model’s performance across various test window sizes denoted by $\Delta$. In this setup, a test window of size $\Delta\times\Delta$ is held out and masked during training, while all other surrounding data points are used for learning. All windows are centered on our default ($Z_{0},N_{0})=(40,62)$ nucleus. Since our graph is built based on spatial nuclear chart connectivity, increasing $\Delta$ reduces the number of available training nodes, making the imputation task increasingly challenging. As shown in Fig. 11, both GraphSAGE and GatedConv exhibit a clear degradation in test PSNR as the window length increases from $\Delta=2$ to $\Delta=8$. This trend aligns with the intuition that accurate imputation demands more effective long-range message passing for larger masked regions. This result highlights the importance of carefully balancing test window size and graph connectivity design. This is especially relevant in nuclear physics applications where the finite extent of the nuclear chart combined with the limited availability of actual data to train poses practical constraints.
@@
Figure 11: Test PSNR values across varying window sizes for GraphSAGE (left) and GatedConv (right) approaches, with the center of the test window fixed at $(Z_{0},N_{0})=(42,60)$. The 2-skip edge connectivity is used for all cases.
VII.3 Loss Space and Decoder Fine-tuning Effects
In this section, building on the training discussion in Section III.3.2, we present an ablation study comparing different optimization strategies for VAE-based embeddings and INR-based embeddings. Specifically, we analyze the effect of minimizing the loss in either the latent space (dimension: $k=32$) or the original space (dimension: $m=256$), with and without decoder fine-tuning. Let $\mathcal{L}_{\text{GNN/32}}$ denote the MSE loss computed between predicted latent vectors and corresponding ground-truth latent vectors, and let $\mathcal{L}_{\text{GNN/256}}$ denote the MSE loss computed between predicted and ground-truth cross sections in the original input space. The result in Table 2 highlights the distinct behaviors between the VAE and INR embeddings under different training regimes. When minimizing the loss in the original input space, the VAE-based embedding achieves the highest PSNR score when the decoder is fine-tuned jointly with the GNN. This suggests that the VAE’s decoder is flexible enough to adapt to latent predictions during training, allowing for better end-to-end optimization. Interestingly, INR embeddings perform best when trained with loss computed directly in the latent space without involving the decoder in the optimization loop. This setup bypasses the complications associated with fine-tuning the INR hypernetwork. These observations justify the different training strategies outlined in Algorithm 1 for each type of embedding.
Table 2: Comparison of test PSNR scores under different training strategies with GatedConv. $\mathcal{L}_{\text{GNN/256}}$: MSE in original cross section space; $\mathcal{L}_{\text{GNN/32}}$: MSE in latent space. Note that for $\mathcal{L}_{\text{GNN/32}}$ (Decoder ON), the decoder does not affect the optimization process as the loss is computed before decoding.
Decoder OFF Decoder ON
Loss Type INR VAE INR VAE
$\mathcal{L}_{\text{GNN/256}}$ 25.85 35.78 34.52 36.40
$\mathcal{L}_{\text{GNN/32}}$ 35.48 33.21
VII.4 Location of latent clusters